### v1

# Atomistic analysis of nematic phase transition in 4-cyano-4-$n$-alkyl biphenyl liquid crystals: Sampling for the first-order phase transition and the free-energy decomposition

Abstract
Molecular dynamics (MD) simulations were conducted using the generalized replica exchange method (gREM) on the 4-cyano-4-$n$-alkylbiphenyl ($n$CB) system with $n=5$, 6, 7, and 8, which exhibits a nematic-isotropic (NI) phase transition. Sampling near the phase transition temperature in systems undergoing first-order phase transitions, such as the NI phase transition, is demanding due to the substantial energy gap between the two phases. To address this, gREM, specifically designed for first-order phase transitions, was utilized to enhance sampling near the NI phase transition temperature. Free-energy calculations based on the energy representation (ER) theory were employed to characterize the NI phase transition. ER evaluates the insertion free energy of $n$CB molecule for both nematic and isotropic phases, revealing a change in the temperature dependence across the NI phase transition. Further decomposition into intermolecular interaction energetic and entropic terms shows quantitatively the balance between these contributions at the NI phase transition temperature.
I Introduction
Liquid crystals are composed of anisotropic molecules known as mesogens, exhibiting various intermediate phases. A typical example is the nematic phase, characterized by long-range orientational order of the molecules without translational order of their centers of mass. The smectic phase, on the other hand, exhibits both orientational and translational order, with the molecules forming a layered structure. Additional intermediate phases include the columnar and cholesteric phases. [1, 2]
The most common mesogens are 4-cyano-$4^{\prime}$-$n$-alkylbiphenyl ($n$CB), which undergoes a nematic-isotropic (NI) phase transition at room temperature. An orientation order parameter is introduced to characterize the nematic phase in relation to the director, which is the orientation axis of the entire system. The orientation order parameter decreases with increasing temperature, typically from 0.6 to 0.4 as the NI phase transition approaches, and drops discontinuously to zero at the transition temperature.
The discontinuous behavior is indicative of a first-order phase transition, as described by the mean-field theory of the NI phase transition. [3, 4, 5] The microscopic models explaining this transition are those of Onsager and Maier–Saupe (MS). The Onsager model is a mean-field model, where the NI transition is derived by the competition between excluded volume and rotational entropy effects between rigid and cylindrical molecules. [6] The MS model is another mean-field model that describes the NI phase transition based on an orientation-dependent interaction as the effective intermolecular potential. [7] The fundamental solutions of the Onsager and MS models are equivalent, as derived by minimizing the free energy with respect to the orientation order parameter at a given temperature.
To obtain a more precise molecular-level understanding beyond mean-field models, it is essential to analyze the free energy of mesogens inserted into both isotropic and nematic phases by varying the temperature across the NI phase transition of $n$CB, particularly through molecular dynamics (MD) simulations. However, this free-energy calculation remains challenging and has not been thoroughly elucidated, despite numerous MD simulations conducted on the NI phase transition of liquid crystal systems. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]
Here, the free energy of mesogens is defined as the insertion free energy associated with transferring a mesogen, considered as a solute molecule, from vacuum into a solution system composed of identical mesogens; the molecules other than the tagged one is the solvent. This is equivalent to the work required to ‘switch on’ the solute-solvent intermolecular interaction. In principle, the free-energy calculation is performed via thermodynamic integration in MD simulations, which requires an ensemble average over specific arrangements of the translational and orientational degrees of freedom between the solute and solvent. In the atomistic description of mesogen molecules, the degrees of freedom exceed six dimensions when accounting for the molecule’s internal degrees of freedom. Thus, a free-energy analysis is a key to understanding an NI phase transition, and in the analysis, it is desirable to address the roles of such intermolecular interactions as electrostatic and van der Waals with keeping the atomic-level resolution.
Figure 1: Temperature dependence of orientational order parameter $S$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The error bar at each temperature corresponds to the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
In this study, we performed MD simulations of $n$CB system with $n=5$, 6, 7, and 8 using the united-atom (UA) model. We also applied the generalized replica-exchange method (gREM), designed to improve sampling of the states near the first-order phase transition by utilizing a non-Boltzmann wight. [30, 31, 32, 33, 34, 35, 36, 37, 38, 39] Note that gREM simulations of 5CB liquid crystals [40] and other applications of the replica-exchange method to coarse-grained liquid crystalline systems were conducted. [41, 42, 43, 44] The insertion free energy was subsequently evaluated using the energy representation (ER) theory for the states obtained through gREM, enabling the assessment of the thermodynamic stability of mesogens in both isotropic and nematic phases across varying temperatures. Within the ER framework, it is also possible to decompose the insertion free energy into the energetic and entropic contributions. We conduct free-energy decomposition and address the roles of these contributions.
Figure 2: (a) Orientational order parameter $S$ as a function of temperature scaled by $T_{\mathrm{NI}}$ from MD simulations for $n$CB ($n=5$, 6, 7, and 8). For comparison, the result of the MS model is also plotted as black dashed curve. (b) Orientation angle distribution function $P(\cos\theta)$ from MD simulations at $T_{\mathrm{NI}}$ for $n$CB ($n=5$, 6, 7, and 8). The corresponding values of the orientation order parameter $S$ are presented. For comparison, the MS model described by Eq. (11) with $S=0.4295$ is also plotted as black dashed curve.
II Model and methods
The molecular model employed was the UA model of the AMBER force field, developed by Tiberio et al. to reproduce the experimental behavior near the NI phase transition temperature. [17] The system consists of $N=4000$ molecules; it was seen in Ref. 40 that the system size with $N=4000$ is enough for treatments of the NI transition. The simulation box is a cubic cell with periodic boundary conditions. All MD simulations in this study were performed using the Large-scale Atomic/Molecular Massively Parallel Simulator (LAMMPS) [45]. The time step was set to 2 fs. Temperature was maintained using Nosé–Hoover thermostat, while pressure was controlled by isotropic Nosé–Hoover barostat with a pressure of 1 atm.
For the isobaric version of gREM, the enthalpy $H(=U+pV)$ dependent non-Boltzmann weight $W_{\alpha}(H)$ $(\alpha=1,\cdots,M)$ is utilized. [34, 35, 36] Here, $U$, $p$, and $V$ represent the potential energy, pressure, and volume of the system, respectively. Here, $\alpha$ denotes the replica index and $M$ is the number of replicas. The weight $w_{\alpha}$ is connected with the effective potential through $w_{\alpha}=-k_{\mathrm{B}}\ln W_{\alpha}$, and is determined by inverse mapping of the effective temperature:
$T_{\alpha}(H)=[{\partial w_{\alpha}}/{\partial H}]^{-1}.$ (1)
The statistical temperature is defined as $T_{\mathrm{S}}(H)=[\partial S/\partial H]^{-1}$, where $S$ is the entropy. At replica $\alpha$, the statistical temperature $T_{\mathrm{S}}$ is evaluated by the most probable value of the enthalpy $H_{\alpha}^{*}$, with $T_{\alpha}(H_{\alpha}^{*})=T_{\mathrm{S}}(H_{\alpha}^{*})$. The simplest expression for $T_{\alpha}$ is the linear effective temperature, given by
$T_{\alpha}(H)=\lambda_{\alpha}+\gamma(H-H_{0}),$ (2)
where $\lambda_{\alpha}$ and $\gamma$ are control parameters representing the intercept and slope of the straight line, respectively, at an arbitrary chosen enthalpy $H_{0}$. Note that $\gamma$ is chosen to be negative so that $T_{\alpha}(H)$ intersects $T_{\mathrm{S}}(H)$ at only one point. In practice, $\gamma$ is selected as
$\gamma=\frac{T_{M}-T_{1}}{\tilde{H}_{1}-\tilde{H}_{M}},$ (3)
where $\hat{H}_{1}$ and $\hat{H}_{M}$ represent the average enthalpies from MD simulations at the predefined minimum and maximum temperatures, $T_{1}$ and $T_{M}$, respectively. Furthermore, $H_{0}$ is set to $\hat{H}_{1}$, and the conditions $\lambda_{1}=T_{1}$ and $\lambda_{M}=T_{M}-\gamma(\hat{H}_{M}-\hat{H}_{1})$ leads to
$\lambda_{\alpha}=\lambda_{1}+(\alpha-1)\Delta\lambda,$ (4)
where $\Delta\lambda=(\lambda_{M}-\lambda_{1})/(M-1)$. The exchange between neighbouring replicas, $\alpha$ (with enthalpy $H$) and $\alpha^{\prime}$ (with enthalpy $H^{\prime}$), is determined by the Metropolis method:
$A_{\alpha,\alpha^{\prime}}=\min[1,\exp(\Delta_{\alpha,\alpha^{\prime}})],$ (5)
where $\Delta_{\alpha,\alpha^{\prime}}=w_{\alpha^{\prime}}(H^{\prime})-w_{\alpha^{\prime}}(H)+w_{\alpha}(H)-w_{\alpha}(H^{\prime})$. In this study, the number of replicas was set to $M=17$. Other parameters, $T_{1}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, $\gamma$ are summarized in Table S1 of the supplementary material. The LAMMPS input files used in this study are open at the repository of Zenodo (see DATA AVAILABILITY statement).
Figure S1 of the supplementary material shows the distribution of the enthalpy normalized by the number of molecules, $H/N$, at each replica, for each replica for $n$CB ($n=5$, 6, 7, and 8). Figure S2 of the supplementary material displays the statistical temperature $T_{\mathrm{S}}$ of each temperature as a function of $H/N$, with the linear effective temperature, as defined in Eq. (2), represented by straight lines. Henceforth, the statistical temperature $T_{\mathrm{S}}$ is referred to simply as $T$.
The ER theory provides the stable and efficient formalism for the free-energy calculation in MD simulations. [46, 47, 48, 49, 50] In the ER theory, a multidimensional coordinate that accounts for the degrees of freedom, including the positions and orientations of solvent and solute molecules, is projected to the intermolecular interaction energy coordinate, $\varepsilon$. Based on the Kirkwood’s charging equation, the expression for the insertion free energy $\Delta\mu$ in the ER theory is given by
[…]
where $u_{\lambda}(\varepsilon)$ represents the potential energy, which continuously varies from 0 to $u$ according to the coupling parameter $\lambda$ $(0\leq\lambda\leq 1)$ such that $u_{0}=0$ and $u_{1}=u$, where $u$ is the potential function of interest between the mesogen molecules. Additionally, $\rho_{\lambda}(\varepsilon)$ denotes the energy distribution function of solvent-solute interaction energy at a given coupling parameter $\lambda$. By integrating by parts, Eq. (6) reduces to
[…]
where the first term represents the average sum of solute-solvent interactions in the target system with $\lambda=1$, and the second term accounts for the free-energy penalty due to the insertion of the solute molecule in Eq. (7). Furthermore, in the ER theory, the second term is approximated by a density functional form $\mathcal{F}$ using $\rho_{0}(\varepsilon)$ with $\lambda=0$, representing the reference solvent system, and $\rho(\varepsilon)$ with $\lambda=1$, representing the target system, as expressed in Eq. (8). In other words, ER theory approximates the insertion free energy from the energy distribution functions of the initial ($\lambda=0$) and final ($\lambda=1$) states, thereby reducing the computational cost by avoiding MD simulations of intermediate states. It has also been reported that the error due to the use of density functional approximation is no greater than the inherent error in the force field. [51] In Eq. (8), the first term is the average interaction energy of the solute with the solvent in the system of interest. The second term corresponds to the free-energy penalty due to the reorganization of the solvent structures with the insertion of the solute. When the solvent degrees of freedom are more restricted by the solute, the second term is more positive. In the following, the first and second terms of Eq. (8) are called energetic and entropic terms, respectively.
Figure 3: Radial distribution function $g(r)$ and orientational correlation functions $G_{1}(r)$ and $G_{2}(r)$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The results are plotted for replicas at $T_{\mathrm{NI}}$, as well as for replicas above and below $T_{\mathrm{NI}}$ (isotropic and nematic phases).
III Results and discussion
III.1 Orientational order
To assess the NI phase transition for $n$CB ($n=5$, 6, 7, and 8), we calculated the temperature dependence of the orientational order parameter, $S$, which is defined as
$S=\frac{1}{N}\sum_{i=1}^{N}P_{2}(\cos\theta_{i})$ (9)
where $\theta_{i}$ represents the orientational angle between the $i$-th molecular axis $\bm{u}_{i}$ and the director axis $\bm{n}$. Furthermore, $P_{2}(x)=(3x^{2}-1)/2$ is the second-order Legendre polynomial. In this study, the molecular axis was defined as the CN bond. The second-rank order parameter tensor,
$\bm{Q}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{3}{2}\bm{u}_{i}\otimes\bm{u}_{i}-\frac{1}{2}\bm{I}\right),$ (10)
is employed to describe the orientational order of biaxial nematic liquid crystals. Here, $\bm{I}$ denotes the identity matrix. Due to the traceless property of $\bm{Q}$, three eigenvalues, $\lambda_{-}$, $\lambda_{0}$, and $\lambda_{+}$ ($\lambda_{-}<\lambda_{0}<\lambda_{+}$), satisfies the condition $\lambda_{-}+\lambda_{0}+\lambda_{+}=0$. The maximum eigenvalue $\lambda_{+}$ coincides with the orientational order parameter $S$, and the corresponding eigenvector represents the director $\bm{n}$. To prevent the order parameter from taking large positive values, particularly in the isotropic phase, $S=-2\lambda_{0}$ was calculated in practice. [52]
The temperature dependence of the ensemble average of S is plotted in Fig. 1. In all liquid crystal system, $S$ exhibits a value near 0 at high temperatures, indicating the isotropic phase. As the temperature decreases a discontinuous behavior is observed at a certain temperature, and $S$ reaches a value around 0.4, signaling a phase transition to the nematic phase. The highest temperature at which $S$ reaches approximately 0.4 was defined as the NI phase transition temperature, $T_{\mathrm{NI}}$. The $S$ value gradually increases as the temperature decreases further, indicating a strengthening of the nematic phase’s orientational order. Although 8CB of the same UA model is known to undergo a smectic phase transition upon further cooling, [17, 19] this study specifically targets the NI phase transition and does not address the smectic-nematic phase transition. Simulated $T_{\mathrm{NI}}$ values for $n$CB ($n=5$, 6, 7, and 8) are 308.7 K (replica 8), 300.0 K (replica 9), 324.3 K (replica 12), and 318.9 K (replica 8), respectively. The results demonstrate that $n$CB exhibits odd-even effects as the number of carbons in the alkyl chain increases, consistent with findings from other MD simulations. [10, 12, 14, 16, 17]
Figure 4: Temperature dependence of the insertion free energy $\Delta\mu$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The error bars at each temperature represent the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
Figure 5: Temperature dependence of the decomposition of $\Delta\mu$ into the average sum of the intermolecular interaction energy, $\langle u\rangle$ (purple) and the entropic contribution $\Delta\mu-\langle u\rangle$ (green) for for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). Note that $\langle u\rangle$ and $\Delta\mu-\langle u\rangle$ are plotted on the left and right $y$-axes, respectively, in units of kcal/mol. The error bars at each temperature represent the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
III.2 Comparison with Maier–Saupe model
It is interesting to compare the MD results with the Onsager and MS models. In these models, the distribution function of the cosine of the orientation angle $\cos\theta$ is assumed to be the form of
$P(\cos\theta)=\frac{1}{Z}\exp[\Gamma_{0}SP_{2}(\cos\theta)],$ (11)
where $\Gamma_{0}$ is the parameter characterizing the degree of the nematic ordering. The normalization factor $Z$ is given by
$Z=2\pi\int_{-1}^{1}\exp[\Gamma_{0}SP_{2}(\cos\theta)]d(\cos\theta),$ (12)
which ensures that the distribution function $P(\cos\theta)$ integrates to 1.
In the Onsager model, $\Gamma_{0}$ is represented by the length $L$ and diameter $D$ of the rigid and cylindrical molecule, as well as the volume fraction $\phi$. This model is considered to be a repulsive model that characterizes the NI phase transition due to the excluded volume effect. In contrast, the MS model incorporates an intermolecular potential between molecules with the potential $V(\cos\theta,S)$ at temperature $T$ given by
$V(\cos\theta,S)=-k_{\mathrm{B}}T\Gamma_{0}SP_{2}(\cos\theta),$ (13)
where $\Gamma_{0}$ characterizes the strength of intermolecular interactions. Equation (11) represents the Boltzmann distribution of $V(\cos\theta,S)$, thereby classifying the MS model as an attractive model.
Note that these two models are common in that the solution is obtained by minimizing the free energy $F$ with respect to $S$. In addition, both the Onsager and MS models have the distribution function of Eq. (11) with a parameter $\Gamma_{0}$. Their difference lies in the origin of $\Gamma_{0}$ and its connection to the molecular properties. In the following, the free energy of the MS model is employed to investigate the temperature dependence of the NI phase transition. The expression of $F$ is given by
$\frac{F}{k_{\mathrm{B}}T}=\frac{1}{2}\Gamma_{0}S^{2}+\ln\left(\frac{Z}{4\pi}\right).$ (14)
The first term can be regarded as the ensemble average of the intermolecular potential, $\langle V(\cos\theta,S)/k_{\mathrm{B}}T\rangle/2=\Gamma_{0}S^{2}/2$, where the factor $1/2$ compensates for the double-counting of interactions. The second term corresponds to the Kullback–Leibler divergence, which quantifies the difference between the distribution $P(\cos\theta)$ and the uniform distribution characterizing the isotropic phase, $1/4\pi$. By using $\Gamma_{0}$ as the control parameter, the orientation order parameter $S$ can be evaluated numerically in the self-consistent manner.
At the NI phase transition, the parameter $\Gamma_{0}\approx 4.54$ yields an orientation order parameter value of $S\approx 0.4295$, which coexists with the trivial solution of $S=0$. Using the corresponding NI phase transition temperature $T_{\mathrm{NI}}$, $S$ as a function of $T/T_{\mathrm{NI}}$ for the MS model is plotted in Fig. 2(a). A comparison with MD simulations for $n$CB ($n=5$, 6, 7, and 8) is also presented, showing overall consistency between MD simulations and MS model. Furthermore, the orientation angle distributions, $P(\cos\theta)$, at $T_{\mathrm{NI}}$ for the MS model is illustrated in Fig. 2(b). The result presents the agreement with MD simulation results for $n$CB, particularly highlighting a significant proportion of molecules aligned parallel to the directer. Slight deviations between MD simulations and the MS model are observed in Fig. 2, which are attributed to the value of $S$ at $T_{\mathrm{NI}}$ being less than the mean-field value of $S=0.4295$.
III.3 Orientatinal correlations
While our focus was on the orientational order parameter, which quantifies the nematic ordering of the entire system, we turn our attention to the orientation correlation functions, $G_{1}(r)$ and $G_{2}(r)$, to gain further insight, particularly regarding short-range ordering. [21, 17, 19] The $G_{1}(r)$ and $G_{2}(r)$ are defined by
$G_{1}(r)$ $=\frac{\langle\delta(r-r_{ij})(\bm{u}_{i}\cdot\bm{u}_{j})\rangle}{\langle\delta(r-r_{ij})\rangle},$ (15)
$G_{2}(r)$ $=\frac{\left\langle\delta(r-r_{ij})[(3/2)(\bm{u}_{i}\cdot\bm{u}_{j})^{2}-(1/2)]\right\rangle}{\langle\delta(r-r_{ij})\rangle},$ (16)
where $\bm{u}_{i}$ is the unit dipole moment in the $i$-th molecule, aligned with the CN bond, as used in the calculation of the orientation order parameter $S$. The distance between the charge centers of the $i$-th and $j$-th molecules is denoted as $r_{ij}$ The functions $G_{1}(r)$ and $G_{2}(r)$ characterize the local polar order and local orientational order between molecules, respectively.
$G_{1}(r)$, $G_{2}(r)$, and the radial distribution function $g(r)$ of the replica at $T_{\mathrm{NI}}$ and its neighboring replicas (isotropic and nematic phases) are presented in Fig. 3. At the distance where $g(r)$ shows a peak, $G_{1}(r)$ indicates a polar-ordered correlation even within the isotropic phase. The presence of oscillation at shorter ranges suggests that neighboring molecules tend to adopt anti-parallel configurations. However, at longer distances, this correlation diminishes in both nematic and isotropic phases, asymptotically approaching zero. Similarly, $G_{2}(r)$ reveals orientation ordering at short distances within the isotropic phase, akin to $G_{1}(r)$. At longer distances, $G_{2}(r)$ converges to finite values in the nematic phase, while in the isotropic phase, the values approach zero, indicating the disappearance of long-range orientation order.
III.4 Energy representation free-energy analysis
Finally, we analyze the thermodynamic stability of the mesogen using the gREM sampling near the NI phase transition temperature. Specifically, we calculated the temperature dependence of the insertion free energy in the $n$CB system based on the ER thory (see Eq. (8)). The results are shown in Fig. 4. Note that $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ were evaluated through MD simulations of systems with $N=4000$ and $N=3999$ molecules, respectively. The results of $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ of the replica at $T_{\mathrm{NI}}$ and its neighboring replicas (isotropic and nematic phases) are presented in Fig. S3 and Fig. S4 of the supplementary material, respectively.
[…]
The average sum of the intermolecular interaction energy, $\langle u\rangle$, decreases with decreasing temperatures and exhibits a discontinuous change at $T_{\mathrm{NI}}$. This indicates that the nematic phase is more stable than the isotropic phase in terms of interaction energy. $\langle u\rangle$ can be further decomposed into Lennard-Jones (LJ) and electrostatic contributions, denoted as $\langle u^{\mathrm{LJ}}\rangle$ and $\langle u^{\mathrm{ele}}\rangle$, respectively. The results are illustrated in Fig. S4 of the supplementary material. It is observed that the stabilization of intermolecular interactions is dominated by the contribution of LJ interactions, as indicated by the large changes in their values compared to those of electrostatic terms. This can be attributed to the fact that the $n$CB mesogen is electrically neutral overall, though it does possess a dipole moment at the CN bond.
[…]
IV Conclusions
In this study, we performed MD simulation of the NI phase transition of the UA model $n$CB ($n=5$, 6, 7, and 8), combined with with gREM. The gREM introducing the linear effective temperature was useful for simulating the NI phase transition, because the thermodynamically unstable states, which originates from the first-order phase transition, can be efficiently sampled. The temperature dependence of the orientation order parameter $S$ was evaluated.
The orientation distribution function calculated from the simulation trajectories is found to be in good agreement with the mean-field MS model. Furthermore, the local orientational order was characterized by the orientation correlation functions, $G_{1}(r)$ and $G_{2}(r)$. In both the nematic and isotropic phases, short-range order was observed, in which the molecules were arranged in antiparallel configurations.
The NI phase transition was described thermodynamically by the free-energy analysis based on the ER theory. The temperature dependence of the insertion free energy $\Delta\mu$ reveals the change in thermodynamic stability associated with the NI phase transition. Notably, the temperature dependence of $\Delta\mu$ changes at $T_{\mathrm{NI}}$, indicating a shift in the temperature dependence of the free energy at this point. Further decomposition into intermolecular interaction energies and entropic terms suggests the competition between these factors, which drives the NI phase transition. These findings are consistent with the MS model, which describe the NI phase transition in terms of the competition between the intermolecular attractive force and entropy. It is important to emphasize that our free-energy analysis based on the ER theory exceeds the mean-field treatment by providing a molecular-level description that incorporates LJ, electrostatic, and entropic contributions.
SUPPLEMENTARY MATERIAL
The supplementary material includes parameters of gREM simulations (Table S1); probability distribution of enthalpy of replicas in gREM simulations (Fig. S1); statistical temperature as a function of enthalpy in gREM simulations (Fig. S2); density profiles of intermolecular interaction energy for ER calculations, $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ (Fig. S3 and Fig. S4); and temperature dependence of LJ and electrostatic energies (Fig. S5).
Acknowledgements.
This work was supported by JSPS KAKENHI Grant-in-Aid Grant Nos. JP24H01719, JP22H04542, JP22K03550, JP23K27313, JP23H02622, JP24K00792, JP19H05718, JP22K18953, JP24K17576, and JP24H01727. Y. I. was also supported by JST ACT-X Grant No. JPMJAX23D3. We are grateful to the Fugaku Supercomputing Project (Nos. JPMXP1020230325 and JPMXP1020230327) and the Data-Driven Material Research Project (No. JPMXP1122714694) from the Ministry of Education, Culture, Sports, Science, and Technology and to Maruho Collaborative Project for Theoretical Pharmaceutics. The numerical calculations were performed at Research Center for Computational Science, Okazaki Research Facilities, National Institutes of Natural Sciences (Project: 24-IMS-C051).
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.
Data availability statement
The LAMMPS input files and initial configurations are available in Zenodo at https://doi.org/10.5281/zenodo.13690896. Further data that support the findings of this study are available from the corresponding author upon reasonable request.
Supplementary Material
Atomistic analysis of nematic phase transition in 4-cyano-4-$n$-alkyl biphenyl liquid crystals: Sampling for the first-order phase transition and the free-energy decomposition
Shunsuke Ogita,1 Yoshiki Ishii,2 Go Watanabe,2 Hitoshi Washizu,3 Kang Kim,1 and Nobuyuki Matubayasi1
1)Division of Chemical Engineering, Graduate School of Engineering Science, Osaka University, Osaka 560-8531, Japan
2)Department of Data Science, School of Frontier Engineering, Kitasato University, Sagamihara, Kanagawa 252-0373, Japan
3)Graduate School of Information Science, University of Hyogo, Hyogo 650-0047, Japan
Table S1: Parameters, $T_{0}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, and $\gamma$, for gREM calculations for $n$CB ($n=5$, 6, 7, and 8).
5CB 6CB 7CB 8CB
$T_{1}$ (K) 300 290 310 310
$T_{M}$ (K) 320 310 330 330
$\tilde{H}_{0}$ (kcal/mol) 275901.8 253970.7 258457.6 278985.2
$\tilde{H}_{M}$ (kcal/mol) 284902.2 262995.8 269201.7 289908
[…]
Figure S1: Probability distributions of enthalpy $H$ of each replica in gREM simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). From left to right, the replica index $\alpha$ ranges from 1 to $M=17$. The enthalpy is normalized by the number of molecules $N=4000$.
Figure S2: Statistical temperature $T_{\mathrm{S}}$ as a function of enthalpy $H$ at each replica in gREM simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). From left to right, the replica index $\alpha$ ranges from 1 to $M=17$. The enthalpy is normalized by the number of molecules $N=4000$. The horizontal color bar indicates the NI transition temperature $T_{\mathrm{NI}}$. The straight line represents the linear effective temperature at each replica, given by $T_{\alpha}(H)=\lambda_{\alpha}+\gamma(H-H_{0})$ (see Eq. (5) in the main text).
Figure S3: Density profiles of intermolecular interaction energy, $\rho(\epsilon)$, obtained from MD simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The results are plotted for replicas at $T_{\mathrm{NI}}$, as well as for replicas above and below $T_{\mathrm{NI}}$ (isotropic and nematic phases).
Figure S4: Density profiles of intermolecular interaction energy, $\rho_{0}(\epsilon)$, obtained from MD simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). The results are plotted for replicas at $T_{\mathrm{NI}}$, as well as for replicas above and below $T_{\mathrm{NI}}$ (isotropic and nematic phases).
Figure S5: Temperature dependence of the decomposition of the intermolecular interaction energy $\langle u\rangle$ into LJ and electrostatic contributions, $\langle u^{\mathrm{LJ}}\rangle$ and $\langle u^{\mathrm{ele}}\rangle$, for 5CB (a), 6CB (b), 7CB (c), and 8CB (d). Note that $\langle u^{\mathrm{LJ}}\rangle$ (purple) and $\langle u^{\mathrm{ele}}\rangle$ (green) are plotted on the left and right $y$-axes, respectively, in units of kcal/mol. The error bars at each temperature represent the standard deviation. The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.

### v2

@@
 However, this free-energy calculation remains challenging and has not been thoroughly elucidated, despite numerous MD simulations conducted on the NI phase transition of liquid crystal systems.
-[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]
+[8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]
 Here, the free energy of mesogens is defined as the insertion free energy associated with transferring a mesogen, considered as a solute molecule, from vacuum into a solution system composed of identical mesogens; the molecules other than the tagged one is the solvent.
@@
 Figure 1: Temperature dependence of orientational order parameter $S$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
+The molecular axis is defined using the CN bond, represented by the purple vector in Fig.
+S1 of the supplementary material.
 The error bar at each temperature corresponds to the standard deviation.
@@
 In this study, we performed MD simulations of $n$CB system with $n=5$, 6, 7, and 8 using the united-atom (UA) model.
-We also applied the generalized replica-exchange method (gREM), designed to improve sampling of the states near the first-order phase transition by utilizing a non-Boltzmann wight.
+We also applied the generalized replica-exchange method (gREM), designed to improve sampling of the states near the first-order phase transition by utilizing a non-Boltzmann weight.
 [31, 32, 33, 34, 35, 36, 37, 38, 39, 40] Note that gREM simulations of 5CB liquid crystals [41] and other applications of the replica-exchange method to coarse-grained liquid crystalline systems were conducted.
@@
 to reproduce the experimental behavior near the NI phase transition temperature.
-[17] The system consists of $N=4000$ molecules; it was seen in Ref.
+[17] The molecular structures of $n$CB ($n=5$, 6, 7, and 8) are illustrated in Fig.
+S1 of the supplementary material.
+The system consists of $N=4000$ molecules; it was seen in Ref.
 41 that the system size with $N=4000$ is enough for treatments of the NI transition.
@@
 The time step was set to 2 fs.
+The total simulation times of gREM were 360 ns for 5CB and 6CB and 450 ns for 7CB and 8CB, respectively, starting from an initial configuration with a random structure.
 Temperature was maintained using Nosé–Hoover thermostat, while pressure was controlled by isotropic Nosé–Hoover barostat with a pressure of 1 atm.
@@
 [35, 36, 37] Here, $U$, $p$, and $V$ represent the potential energy, pressure, and volume of the system, respectively.
-Here, $\alpha$ denotes the replica index and $M$ is the number of replicas.
+In addition, $\alpha$ denotes the replica index and $M$ is the number of replicas.
 The weight $w_{\alpha}$ is connected with the effective potential through $w_{\alpha}=-k_{\mathrm{B}}\ln W_{\alpha}$, and is determined by inverse mapping of the effective temperature:
 $T_{\alpha}(H)=[{\partial w_{\alpha}}/{\partial H}]^{-1}.$ (1)
-The statistical temperature is defined as $T_{\mathrm{S}}(H)=[\partial S/\partial H]^{-1}$, where $S$ is the entropy.
-At replica $\alpha$, the statistical temperature $T_{\mathrm{S}}$ is evaluated by the most probable value of the enthalpy $H_{\alpha}^{*}$, with $T_{\alpha}(H_{\alpha}^{*})=T_{\mathrm{S}}(H_{\alpha}^{*})$.
+The statistical temperature is defined as $T_{\mathcal{S}}(H)=[\partial\mathcal{S}/\partial H]^{-1}$, where $\mathcal{S}$ is the entropy.
+At replica $\alpha$, the statistical temperature $T_{\mathcal{S}}$ is evaluated by the most probable value of the enthalpy $H_{\alpha}^{*}$, with $T_{\alpha}(H_{\alpha}^{*})=T_{\mathcal{S}}(H_{\alpha}^{*})$.
 The simplest expression for $T_{\alpha}$ is the linear effective temperature, given by
@@
 where $\lambda_{\alpha}$ and $\gamma$ are control parameters representing the intercept and slope of the straight line, respectively, at an arbitrary chosen enthalpy $H_{0}$.
-Note that $\gamma$ is chosen to be negative so that $T_{\alpha}(H)$ intersects $T_{\mathrm{S}}(H)$ at only one point.
+Note that $\gamma$ is chosen to be negative so that $T_{\alpha}(H)$ intersects $T_{\mathcal{S}}(H)$ at only one point.
 In practice, $\gamma$ is selected as
@@
 Other parameters, $T_{1}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, $\gamma$ are summarized in Table S1 of the supplementary material.
-The LAMMPS input files used in this study are open at the repository of Zenodo (see DATA AVAILABILITY statement).
+The LAMMPS input files used in this study and 1 ns trajectory files are openly available at the repository of Zenodo (see DATA AVAILABILITY statement).
 Figure S2 of the supplementary material shows the distribution of the enthalpy normalized by the number of molecules, $H/N$, at each replica, for each replica for $n$CB ($n=5$, 6, 7, and 8).
-Figure S2 of the supplementary material displays the statistical temperature $T_{\mathrm{S}}$ of each temperature as a function of $H/N$, with the linear effective temperature, as defined in Eq.
+These results were calculated from the final 36 ns (5CB and 6CB) and 108 ns (7CB and 8CB) trajectories in the gREM calculation.
+The findings demonstrate sufficient overlap of enthalpy distributions between replicas.
+However, the distributions are broad in certain regions of enthalpy between specific replicas, which can be attributed to enthalpy gaps caused by the NI phase transition.
+Despite this, we confirmed that the replica exchange acceptance ratio remains maintained, as shown in Fig.
+S3 of the supplementary material.
+To enhance the replica exchange acceptance ratio, particularly for larger system sizes, it is necessary to increase the number of replicas $M$, significantly raising computational demands.
+Figure S4 of the supplementary material displays the statistical temperature $T_{\mathcal{S}}$ of each temperature as a function of $H/N$, with the linear effective temperature, as defined in Eq.
 (2), represented by straight lines.
-Henceforth, the statistical temperature $T_{\mathrm{S}}$ is referred to simply as $T$.
+Henceforth, the statistical temperature $T_{\mathcal{S}}$ is referred to simply as $T$.
 The ER theory provides the stable and efficient formalism for the free-energy calculation in MD simulations.
@@
 It has also been reported that the error due to the use of density functional approximation is no greater than the inherent error in the force field.
+[52] Specifically, the ER method was used to calculate the solvation free energy of amino acid analogs in a pure-water solvent, showing a typical discrepancy of less than 1 kcal/mol compared to the numerically exact values obtained using the Bennett acceptance ratio (BAR) method.
 [52] In Eq.
@@
 $S=\frac{1}{N}\sum_{i=1}^{N}P_{2}(\cos\theta_{i})$ (9)
-where $\theta_{i}$ represents the orientational angle between the $i$-th molecular axis $\bm{u}_{i}$ and the director axis $\bm{n}$.
+where $\theta_{i}$ represents the orientational angle between the $i$-th molecular axis $\bm{u}_{i}$ and the director $\bm{n}$.
 Furthermore, $P_{2}(x)=(3x^{2}-1)/2$ is the second-order Legendre polynomial.
-In this study, the molecular axis was defined as the CN bond.
-The second-rank order parameter tensor,
+In MD simulations, the direction of the director $\bm{n}$ cannot be predetermined. Alternatively, the second-rank order parameter tensor,
 $\bm{Q}=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{3}{2}\bm{u}_{i}\otimes\bm{u}_{i}-\frac{1}{2}\bm{I}\right),$ (10)
-is employed to describe the orientational order of biaxial nematic liquid crystals.
+is employed to analyze the orientational order of biaxial nematic liquid crystals.
 Here, $\bm{I}$ denotes the identity matrix.
 Due to the traceless property of $\bm{Q}$, three eigenvalues, $\lambda_{-}$, $\lambda_{0}$, and $\lambda_{+}$ ($\lambda_{-}<\lambda_{0}<\lambda_{+}$), satisfies the condition $\lambda_{-}+\lambda_{0}+\lambda_{+}=0$.
-The maximum eigenvalue $\lambda_{+}$ coincides with the orientational order parameter $S$, and the corresponding eigenvector represents the director $\bm{n}$.
-To prevent the order parameter from taking large positive values, particularly in the isotropic phase, $S=-2\lambda_{0}$ was calculated in practice.
-[52]
+The orientational order parameter $S$ can be evaluated by the maximum eigenvalue $\lambda_{+}$ as $S=\lambda_{+}$, and the corresponding eigenvector represents the director $\bm{n}$.
+However, this definition results in positive values for $S$ even in the isotropic phase due to the traceless property of $Q$.
+In this study, following the approach of Eppenga and Frenkel, $S$ was calculated by $-2\lambda_{0}$, considering $\lambda_{0}\approx\lambda_{-}=-\lambda_{+}/2$ in the nematic phase.
+[53] This definition ensures that $S$ fluctuates around zero in the isotropic phase, indicating the absence of orientational order.
 The temperature dependence of the ensemble average of $S$ is plotted in Fig.
 1.
-In all liquid crystal system, $S$ exhibits a value near 0 at high temperatures, indicating the isotropic phase.
-As the temperature decreases a discontinuous behavior is observed at a certain temperature, and $S$ reaches a value around 0.4, signaling a phase transition to the nematic phase.
+In this study, the molecular axis was defined as the CN bond, represented by the purple vector in Fig.
+S1 of the supplementary material.
+Here, $S$ was calculated from the final 36 ns (5CB and 6CB) and 108 ns (7CB and 8CB) trajectories in the gREM simulation, consistent with the calculation of the enthalpy distribution shown in Fig.
+S2 of the supplementary material.
+In all liquid crystal systems, $S$ exhibits a value near 0 at high temperatures, indicating the isotropic phase.
+As the temperature decreases, a discontinuous behavior is observed at a certain temperature, and $S$ reaches a value around 0.4, signaling a phase transition to the nematic phase.
 The highest temperature at which $S$ reaches approximately 0.4 was defined as the NI phase transition temperature, $T_{\mathrm{NI}}$.
@@
 The results demonstrate that $n$CB exhibits odd-even effects as the number of carbons in the alkyl chain increases, consistent with findings from other MD simulations.
-[10, 12, 14, 16, 17]
+[10, 12, 14, 16, 17] We also calculated the orientational order parameter $S$ using an alternative definition of the molecular axis, defined by the line connecting N and C of the phenyl ring, represented by the red vector in Fig.
+S1 of the supplementary material.
+This definition was examined in a previous study.
+[30] The temperature dependence of $S$, shown in Fig.
+S5 of the supplementary material, demonstrates that $S$ remains unaffected by the choice of the molecular axis.
 Figure 4: Temperature dependence of the insertion free energy $\Delta\mu$ for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
@@
 Equation (11) represents the Boltzmann distribution of $V(\cos\theta,S)$, thereby classifying the MS model as an attractive model.
+In this context, the MS model considers the combined effects of volume fraction and temperature on the NI phase transition, making it more suitable for describing thermotropic liquid crystals compared to the Onsager model.
 Note that these two models are common in that the solution is obtained by minimizing the free energy $F$ with respect to $S$.
@@
 The second term corresponds to the Kullback–Leibler divergence, which quantifies the difference between the distribution $P(\cos\theta)$ and the uniform distribution characterizing the isotropic phase, $1/4\pi$.
-By using $\Gamma_{0}$ as the control parameter, the orientation order parameter $S$ can be evaluated numerically in the self-consistent manner.
+By using $\Gamma_{0}$ as the control parameter, the orientational order parameter $S$ can be evaluated numerically in the self-consistent manner based on Eq.
+(14).
 At the NI phase transition, the parameter $\Gamma_{0}\approx 4.54$ yields an orientation order parameter value of $S\approx 0.4295$, which coexists with the trivial solution of $S=0$.
@@
 2(b).
-The result presents the agreement with MD simulation results for $n$CB, particularly highlighting a significant proportion of molecules aligned parallel to the directer.
+The result presents the agreement with MD simulation results for $n$CB, particularly highlighting a significant proportion of molecules aligned parallel to the director.
 Slight deviations between MD simulations and the MS model are observed in Fig.
@@
 III.4 Energy representation free-energy analysis
-Finally, we analyze the thermodynamic stability of the mesogen using the gREM sampling near the NI phase transition temperature.
-Specifically, we calculated the temperature dependence of the insertion free energy in the $n$CB system based on the ER thory (see Eq.
+We analyze the thermodynamic stability of the mesogen using the gREM sampling near the NI phase transition temperature.
+Specifically, we calculated the temperature dependence of the insertion free energy, $\Delta\mu$, in the $n$CB system based on the ER thory (see Eq.
 (8)).
@@
 The results of $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ of the replica at $T_{\mathrm{NI}}$ and its neighboring replicas (isotropic and nematic phases) are presented in Fig.
-S3 and Fig.
-S4 of the supplementary material, respectively.
+S6 and Fig.
+S7 of the supplementary material, respectively.
 […]
@@
 The results are illustrated in Fig.
-S4 of the supplementary material.
+S8 of the supplementary material.
 It is observed that the stabilization of intermolecular interactions is dominated by the contribution of LJ interactions, as indicated by the large changes in their values compared to those of electrostatic terms.
@@
 […]
+Finally, we discuss the odd-even effect based on our free-energy calculation of Fig.
+5.
+For odd-numbered alkyl chains, the terminal alkyl group aligns parallel to the molecular axis, while for even-numbered alkyl chains, the angle to the molecular axis tends to be larger.
+This results in reduced molecular anisotropy in the even-numbered compounds, leading to a lower NI phase transition temperature.
+We analyzed the energy gap and entropy gap across the NI transition temperature, denoted as $\Delta\langle u\rangle$ and $\Delta(\Delta\mu-\langle u\rangle)$.
+As illustrated in Fig.
+S9 of the supplementary material, the results indicate that odd-numbered compounds exhibit more stability of the nematic ordering with a significantly large energy loss, accompanied by a large entropy loss.
+This suggests that the nematic phase in odd-numbered compounds is stabilized by a strong molecular anisotropy.
+Conversely, the even-numbered compounds show smaller energy and entropy gaps, reflecting their reduced anisotropy and lower transition temperature to the nematic phase.
 IV Conclusions
@@
 SUPPLEMENTARY MATERIAL
-The supplementary material includes parameters of gREM simulations (Table S1); probability distribution of enthalpy of replicas in gREM simulations (Fig.
-S1); statistical temperature as a function of enthalpy in gREM simulations (Fig.
-S2); density profiles of intermolecular interaction energy for ER calculations, $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ (Fig.
-S3 and Fig.
-S4); and temperature dependence of LJ and electrostatic energies (Fig.
-S5).
+The supplementary material includes parameters of gREM simulations (Table S1); molecular structures of $n$CB (Fig.
+S1); probability distribution of enthalpy of replicas in gREM simulations (Fig.
+S2); acceptance ratios of replica exchanges (Fig.
+S3); statistical temperature as a function of enthalpy in gREM simulations (Fig.
+S4); orientational order parameter $S$ using another molecular axis (Fig.
+S5); density profiles of intermolecular interaction energy for ER calculations, $\rho(\varepsilon)$ and $\rho_{0}(\varepsilon)$ (Fig.
+S6 and Fig.
+S7); and temperature dependence of LJ and electrostatic energies (Fig.
+S8); dependence of energy gap $\Delta\langle u\rangle$ and entropy gap $\Delta(\Delta\mu-\langle u\rangle)$ on the number of carbons in the alkyl chain (Fig.
+S9).
 Acknowledgements.
@@
 Data availability statement
-The LAMMPS input files and initial configurations are available in Zenodo at https://doi.org/10.5281/zenodo.13690896.
+The LAMMPS input files and initial configurations and 1 ns trajectories are available in Zenodo at https://doi.org/10.5281/zenodo.13690896.
 Further data that support the findings of this study are available from the corresponding author upon reasonable request.
-Supplementary Material
-Atomistic analysis of nematic phase transition in 4-cyano-4-$n$-alkyl biphenyl liquid crystals: Sampling for the first-order phase transition and the free-energy decomposition
-Shunsuke Ogita,1 Yoshiki Ishii,2 Go Watanabe,2 Hitoshi Washizu,3 Kang Kim,1 and Nobuyuki Matubayasi1
-1)Division of Chemical Engineering, Graduate School of Engineering Science, Osaka University, Osaka 560-8531, Japan
-2)Department of Data Science, School of Frontier Engineering, Kitasato University, Sagamihara, Kanagawa 252-0373, Japan
-3)Graduate School of Information Science, University of Hyogo, Hyogo 650-0047, Japan
-Table S1: Parameters, $T_{0}$, $T_{M}$, $\hat{H}_{1}$, $\hat{H}_{M}$, and $\gamma$, for gREM calculations for $n$CB ($n=5$, 6, 7, and 8).
-5CB 6CB 7CB 8CB
-$T_{1}$ (K) 300 290 310 310
-$T_{M}$ (K) 320 310 330 330
-$\tilde{H}_{0}$ (kcal/mol) 275901.8 253970.7 258457.6 278985.2
-$\tilde{H}_{M}$ (kcal/mol) 284902.2 262995.8 269201.7 289908
-[…]
-Figure S1: Probability distributions of enthalpy $H$ of each replica in gREM simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
-From left to right, the replica index $\alpha$ ranges from 1 to $M=17$.
-The enthalpy is normalized by the number of molecules $N=4000$.
-Figure S2: Statistical temperature $T_{\mathrm{S}}$ as a function of enthalpy $H$ at each replica in gREM simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
-From left to right, the replica index $\alpha$ ranges from 1 to $M=17$.
-The enthalpy is normalized by the number of molecules $N=4000$.
-The horizontal color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.
-The straight line represents the linear effective temperature at each replica, given by $T_{\alpha}(H)=\lambda_{\alpha}+\gamma(H-H_{0})$ (see Eq.
-(5) in the main text).
-Figure S3: Density profiles of intermolecular interaction energy, $\rho(\epsilon)$, obtained from MD simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
-The results are plotted for replicas at $T_{\mathrm{NI}}$, as well as for replicas above and below $T_{\mathrm{NI}}$ (isotropic and nematic phases).
-Figure S4: Density profiles of intermolecular interaction energy, $\rho_{0}(\epsilon)$, obtained from MD simulations for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
-The results are plotted for replicas at $T_{\mathrm{NI}}$, as well as for replicas above and below $T_{\mathrm{NI}}$ (isotropic and nematic phases).
-Figure S5: Temperature dependence of the decomposition of the intermolecular interaction energy $\langle u\rangle$ into LJ and electrostatic contributions, $\langle u^{\mathrm{LJ}}\rangle$ and $\langle u^{\mathrm{ele}}\rangle$, for 5CB (a), 6CB (b), 7CB (c), and 8CB (d).
-Note that $\langle u^{\mathrm{LJ}}\rangle$ (purple) and $\langle u^{\mathrm{ele}}\rangle$ (green) are plotted on the left and right $y$-axes, respectively, in units of kcal/mol.
-The error bars at each temperature represent the standard deviation.
-The vertical color bar indicates the NI transition temperature $T_{\mathrm{NI}}$.